### v2

@@ 1 Introduction
Medical image segmentation is a crucial for disease diagnosis and quantitative assessment across the entire clinical workflow [Antonelli et al.(2022)Antonelli, Reinke, Bakas, Farahani, Kopp-Schneider, Landman, Litjens, Menze, Ronneberger, Summers, et al.]. This technology is especially critical in biomedical applications, where detecting individual tumors, organs, or cell entities and segmenting their shapes represents a major bottleneck [Fiaz et al.(2022)Fiaz, Heidari, Anwer, and Cholakkal]. The inherent challenges of automated segmentation arise from the considerable variations in the size, shape, appearance, and density of the object of interest. Deep learning networks excel in medical image analysis, outperforming traditional SOTA techniques. The fully convolutional neural network (FCN) stands out as one of the pioneering approaches employed for image segmentation [Long et al.(2015)Long, Shelhamer, and Darrell]. The seminal U-Net [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox], featuring a series of convolutional and down-sampling layers to gather contextual information along its contracting path, with skip connections from the encoder to up-sample both coarse and fine features, has demonstrated eminent segmentation potential. The CNN and U-Net architectures have been enhanced through various adaptations aimed at optimizing their performance. Innovations have included the introduction of attention mechanisms [Oktay et al.(2018)Oktay, Schlemper, Folgoc, Lee, Heinrich, Misawa, Mori, McDonagh, Hammerla, Kainz, et al.], the incorporation of dynamic modeling [Shen et al.(2023)Shen, Wang, Zhang, Qin, Jin, and Zhao, Li et al.(2019b)Li, Wang, Hu, and Yang, Azad et al.(2023b)Azad, Niggemeier, Huttemann, Kazerouni, Aghdam, Velichko, Bagci, and Merhof], modifying skip connections [Huang et al.(2020)Huang, Lin, Tong, Hu, Zhang, Iwamoto, Han, Chen, and Wu], and the replacement of backbone modules within segmentation networks [Karaali et al.(2022)Karaali, Dahyot, and Sexton]. These enhancements have been critical in advancing the capabilities of these architectures to handle the complexities of medical image segmentation. While CNNs have traditionally been preferred for this task, they struggle to effectively model long-range dependencies and are limited by static receptive fields, which restrict their ability to capture the shape and structural details needed for precise medical image segmentation. The Vision Transformer (ViT) [Dosovitskiy et al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et al.] was designed to address this limitation. Equipped with the self-attention mechanism, ViTs excel at capturing global information and modeling long-range dependencies. However, transformers are limited in capturing the local representation and context and are computationally inefficient with quadratic complexity. Several methods have been proposed to overcome the disadvantages of the transformer model [Khan et al.(2022)Khan, Naseer, Hayat, Zamir, Khan, and Shah], such as hybrid CNN-Transformer architectures [Heidari et al.(2023)Heidari, Kazerouni, Soltany, Azad, Aghdam, Cohen-Adad, and Merhof, Sadegheih et al.(2024)Sadegheih, Bozorgpour, Kumari, Azad, and Merhof], reducing the computational cost of self-attention [Shen et al.(2024)Shen, Zhang, Zhao, Yi, and Li], performing the self-attention mechanism in local regions [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], or calculating channel attention instead of spatial attention [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.]. However, these methods capture only partially the global context while addressing the diverse challenges in medical images, such as variations in form, visual characteristics, and compactness, necessitating the integration of both local and global features. A feasible approach in addressing the inherent limitations of the aforementioned architectures with respect to local and global information processing is the use of skip connections [Azad et al.(2022)Azad, Aghdam, Rauland, Jia, Avval, Bozorgpour, Karimijafarbigloo, Cohen, Adeli, and Merhof]. In medical image segmentation, capturing both local details (like the texture of tissues) and global context (such as the structure of organs) is crucial. Skip connections help in preserving high-resolution details through the network layers, enabling the model to use both fine and coarse features for more precise segmentation. Skip connections assist CNNs by reintegrating precise local features from initial layers into deeper ones, blending detailed and broad information for comprehensive contextual understanding [Dai et al.(2020)Dai, Gieseke, Oehmcke, Wu, and Barnard], thus overcoming their naturally limited receptive fields. For transformers, skip connections enhance performance by integrating local details directly into layers focused on global processing, effectively balancing local precision with global awareness. However, most of these studies have focused solely on improving the encoder or decoder, and not on altering the original skip connection design as such [Wang et al.(2022a)Wang, Cao, Wang, and Zaiane]. Therefore, we propose MSANet, augmented by innovative skip connections that incorporate our novel Multi-Scale Adaptive Spatial Attention Gate (MASAG). This module strategically highlights spatially relevant features across multiple scales, enabling it to effectively model complex structures for delineating and localizing objects of interest. This stage facilitates a spatial interplay between local and global features, enriching the feature maps with a nuanced, contextually informed representation [Fan et al.(2023)Fan, Huang, Zhou, and He]. Within the decoder, initial layers employ the Large Kernel Attention (LKA) module [Guo et al.(2022)Guo, Lu, Liu, Cheng, and Hu] for adept local and global feature management, addressing typical high-resolution image processing challenges in transformers. Deeper layers of the decoder use Dual Attention Enhanced Transformer blocks (DAE-Former) [Azad et al.(2023a)Azad, Arimond, Aghdam, Kazerouni, and Merhof], effectively preserving long-range interactions in lower-resolution images while incorporating both spatial and channel attentions seamlessly, thus maintaining the grid integrity of the images. This strategy guarantees accurate and context-sensitive delineation of anatomical features within intricate medical images.
@@ 2 Related Work
Transformer-based and Hybrid Segmentation Networks: Transformers have demonstrated their capability to gather global information, addressing the drawback of CNNs in capturing long-range semantic dependencies and limited receptive fields. Some of these studies aim to capture long-range spatial context using purely transformer-based approaches instead of CNN-based methods, which partially broaden the receptive field of CNNs [Valanarasu et al.(2021)Valanarasu, Oza, Hacihaliloglu, and Patel]. However, the quadratic computational complexity of transformers, along with their inefficiency in capturing local contextual relationships between intra-path pixels, limits their usability in dense vision tasks such as medical image segmentation, which requires the neighboring pixel dependencies to be considered in the multi-scale and hierarchical pattern [Khan et al.(2023)Khan, Rauf, Khan, Rathore, Khan, Shah, Farooq, Asif, Asif, Zahoora, Khalil, Qamar, Asif, Khan, Majid, and Gwak]. To concurrently exploit the multi-scale representation along with the ability to capture local semantic and texture information of CNNs, and long-range interaction of transformers, cohort studies are dedicated to hybrid CNN-transformer based methods [Chen et al.(2021)Chen, Lu, Yu, Luo, Adeli, Wang, Lu, Yuille, and Zhou, Sadegheih et al.(2024)Sadegheih, Bozorgpour, Kumari, Azad, and Merhof, Azad et al.(2023b)Azad, Niggemeier, Huttemann, Kazerouni, Aghdam, Velichko, Bagci, and Merhof]. Specifically, Swin-UNet [Cao et al.(2022)Cao, Wang, Chen, Jiang, Zhang, Tian, and Wang] and DS-TransUNet [Lin et al.(2022)Lin, Chen, Xu, Zhang, Lu, and Zhang] introduce models with a U-shaped structure that utilizes the Swin Transformer [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] for 2D image segmentation.
@@ 3.2 Overall Architecture
The overall architecture of our novel method, named MSANet, as illustrated in Figure 1. The proposed method leverages the MaxViT [Tu et al.(2022)Tu, Talebi, Zhang, Yang, Milanfar, Bovik, and Li] hierarchical transformer in the encoder to effectively capture multi-scale features through a combination of MBConv [Sandler et al.(2019)Sandler, Howard, Zhu, Zhmoginov, and Chen] and various attention mechanisms. Enhanced by innovative skip connections that include our novel MASAG module, the method ensures that spatially relevant features are selectively emphasized while background distractions are minimized due to dynamically adjust of the spatial receptive fields, guided by inputs from both the encoder and decoder. Moreover, MASAG introduces a spatial interaction stage, enabling a bidirectional process between local and global features, thus providing the feature maps with a detailed, contextually informed representation. The decoder’s shallow layers employ the LKA [Guo et al.(2022)Guo, Lu, Liu, Cheng, and Hu] modules to adeptly balance local and global features as well as channel interactions, overcoming typical high-resolution image processing challenges associated with self-attention. Deeper decoder layers utilize DAE-Former blocks [Azad et al.(2023a)Azad, Arimond, Aghdam, Kazerouni, and Merhof] to efficiently maintain long-range dependencies in lower-resolution images, integrating spatial and channel attentions without disrupting the 2D structure of images. Overall, this method aims at accurate, context-aware segmentation of anatomical structures in complex medical images.
@@ 4.3 Quantitative Results
| ScaleFormer [Huang et al.(2022)Huang, Xie1, Lin, Iwamoto, Han, Chen, and Tong] | 111.6 | 48.93 | 89.40 | 83.31 | 86.36 | 74.97 | 95.12 | 80.14 | 88.73 | 64.85 | 82.86 | 16.81 |
@@
| DAEFormer [Azad et al.(2023a)Azad, Arimond, Aghdam, Kazerouni, and Merhof] | 48.07 | 27.89 | 91.82 | 82.39 | 87.66 | 71.65 | 95.08 | 80.77 | 87.84 | 63.93 | 82.63 | 16.39 |
@@
| 2D D-LKA Net [Azad et al.(2023b)Azad, Niggemeier, Huttemann, Kazerouni, Aghdam, Velichko, Bagci, and Merhof] | 101.64 | 19.92 | 91.22 | 84.92 | 88.38 | 73.79 | 94.88 | 84.94 | 88.34 | 67.71 | 84.27 | 20.04 |

### v3

@@ 1 Introduction
Medical image segmentation is a crucial for disease diagnosis and quantitative assessment across the entire clinical workflow [Antonelli et al.(2022)Antonelli, Reinke, Bakas, Farahani, Kopp-Schneider, Landman, Litjens, Menze, Ronneberger, Summers, et al.]. This technology is especially critical in biomedical applications, where detecting individual tumors, organs, or cell entities and segmenting their shapes represents a major bottleneck [Fiaz et al.(2022)Fiaz, Heidari, Anwer, and Cholakkal]. The inherent challenges of automated segmentation arise from the considerable variations in the size, shape, appearance, and density of the object of interest. Deep learning networks excel in medical image analysis, outperforming traditional SOTA techniques. The fully convolutional neural network (FCN) stands out as one of the pioneering approaches employed for image segmentation [Long et al.(2015)Long, Shelhamer, and Darrell]. The seminal U-Net [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox], featuring a series of convolutional and down-sampling layers to gather contextual information along its contracting path, with skip connections from the encoder to up-sample both coarse and fine features, has demonstrated eminent segmentation potential. The CNN and U-Net architectures have been enhanced through various adaptations aimed at optimizing their performance. Innovations have included the introduction of attention mechanisms [Oktay et al.(2018)Oktay, Schlemper, Folgoc, Lee, Heinrich, Misawa, Mori, McDonagh, Hammerla, Kainz, et al.], the incorporation of dynamic modeling [Shen et al.(2023)Shen, Wang, Zhang, Qin, Jin, and Zhao, Li et al.(2019b)Li, Wang, Hu, and Yang, Azad et al.(2024)Azad, Niggemeier, Hüttemann, Kazerouni, Aghdam, Velichko, Bagci, and Merhof], modifying skip connections [Huang et al.(2020)Huang, Lin, Tong, Hu, Zhang, Iwamoto, Han, Chen, and Wu], and the replacement of backbone modules within segmentation networks [Karaali et al.(2022)Karaali, Dahyot, and Sexton]. These enhancements have been critical in advancing the capabilities of these architectures to handle the complexities of medical image segmentation. While CNNs have traditionally been preferred for this task, they struggle to effectively model long-range dependencies and are limited by static receptive fields, which restrict their ability to capture the shape and structural details needed for precise medical image segmentation. The Vision Transformer (ViT) [Dosovitskiy et al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et al.] was designed to address this limitation. Equipped with the self-attention mechanism, ViTs excel at capturing global information and modeling long-range dependencies. However, transformers are limited in capturing the local representation and context and are computationally inefficient with quadratic complexity. Several methods have been proposed to overcome the disadvantages of the transformer model [Khan et al.(2022)Khan, Naseer, Hayat, Zamir, Khan, and Shah], such as hybrid CNN-Transformer architectures [Heidari et al.(2023)Heidari, Kazerouni, Soltany, Azad, Aghdam, Cohen-Adad, and Merhof, Sadegheih et al.(2024)Sadegheih, Bozorgpour, Kumari, Azad, and Merhof], reducing the computational cost of self-attention [Shen et al.(2021)Shen, Zhang, Zhao, Yi, and Li], performing the self-attention mechanism in local regions [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], or calculating channel attention instead of spatial attention [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.]. However, these methods capture only partially the global context while addressing the diverse challenges in medical images, such as variations in form, visual characteristics, and compactness, necessitating the integration of both local and global features. A feasible approach in addressing the inherent limitations of the aforementioned architectures with respect to local and global information processing is the use of skip connections [Azad et al.(2022)Azad, Aghdam, Rauland, Jia, Avval, Bozorgpour, Karimijafarbigloo, Cohen, Adeli, and Merhof]. In medical image segmentation, capturing both local details (like the texture of tissues) and global context (such as the structure of organs) is crucial. Skip connections help in preserving high-resolution details through the network layers, enabling the model to use both fine and coarse features for more precise segmentation. Skip connections assist CNNs by reintegrating precise local features from initial layers into deeper ones, blending detailed and broad information for comprehensive contextual understanding [Dai et al.(2021)Dai, Gieseke, Oehmcke, Wu, and Barnard], thus overcoming their naturally limited receptive fields. For transformers, skip connections enhance performance by integrating local details directly into layers focused on global processing, effectively balancing local precision with global awareness. However, most of these studies have focused solely on improving the encoder or decoder, and not on altering the original skip connection design as such [Wang et al.(2022a)Wang, Cao, Wang, and Zaiane]. Therefore, we propose MSANet, augmented by innovative skip connections that incorporate our novel Multi-Scale Adaptive Spatial Attention Gate (MASAG). This module strategically highlights spatially relevant features across multiple scales, enabling it to effectively model complex structures for delineating and localizing objects of interest. This stage facilitates a spatial interplay between local and global features, enriching the feature maps with a nuanced, contextually informed representation [Fan et al.(2024)Fan, Huang, Zhou, and He]. Within the decoder, initial layers employ the Large Kernel Attention (LKA) module [Guo et al.(2023)Guo, Lu, Liu, Cheng, and Hu] for adept local and global feature management, addressing typical high-resolution image processing challenges in transformers. Deeper layers of the decoder use Dual Attention Enhanced Transformer blocks (DAE-Former) [Azad et al.(2023)Azad, Arimond, Aghdam, Kazerouni, and Merhof], effectively preserving long-range interactions in lower-resolution images while incorporating both spatial and channel attentions seamlessly, thus maintaining the grid integrity of the images. This strategy guarantees accurate and context-sensitive delineation of anatomical features within intricate medical images.
@@ 2 Related Work
Transformer-based and Hybrid Segmentation Networks: Transformers have demonstrated their capability to gather global information, addressing the drawback of CNNs in capturing long-range semantic dependencies and limited receptive fields. Some of these studies aim to capture long-range spatial context using purely transformer-based approaches instead of CNN-based methods, which partially broaden the receptive field of CNNs [Valanarasu et al.(2021)Valanarasu, Oza, Hacihaliloglu, and Patel]. However, the quadratic computational complexity of transformers, along with their inefficiency in capturing local contextual relationships between intra-path pixels, limits their usability in dense vision tasks such as medical image segmentation, which requires the neighboring pixel dependencies to be considered in the multi-scale and hierarchical pattern [Khan et al.(2023)Khan, Rauf, Khan, Rathore, Khan, Shah, Farooq, Asif, Asif, Zahoora, et al.]. To concurrently exploit the multi-scale representation along with the ability to capture local semantic and texture information of CNNs, and long-range interaction of transformers, cohort studies are dedicated to hybrid CNN-transformer based methods [Chen et al.(2021)Chen, Lu, Yu, Luo, Adeli, Wang, Lu, Yuille, and Zhou, Sadegheih et al.(2024)Sadegheih, Bozorgpour, Kumari, Azad, and Merhof, Azad et al.(2024)Azad, Niggemeier, Hüttemann, Kazerouni, Aghdam, Velichko, Bagci, and Merhof]. Specifically, Swin-UNet [Cao et al.(2022)Cao, Wang, Chen, Jiang, Zhang, Tian, and Wang] and DS-TransUNet [Lin et al.(2022)Lin, Chen, Xu, Zhang, Lu, and Zhang] introduce models with a U-shaped structure that utilizes the Swin Transformer [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] for 2D image segmentation.
@@ 3.2 Overall Architecture
The overall architecture of our novel method, named MSANet, as illustrated in Figure 1. The proposed method leverages the MaxViT [Tu et al.(2022)Tu, Talebi, Zhang, Yang, Milanfar, Bovik, and Li] hierarchical transformer in the encoder to effectively capture multi-scale features through a combination of MBConv [Sandler et al.(2018)Sandler, Howard, Zhu, Zhmoginov, and Chen] and various attention mechanisms. Enhanced by innovative skip connections that include our novel MASAG module, the method ensures that spatially relevant features are selectively emphasized while background distractions are minimized due to dynamically adjust of the spatial receptive fields, guided by inputs from both the encoder and decoder. Moreover, MASAG introduces a spatial interaction stage, enabling a bidirectional process between local and global features, thus providing the feature maps with a detailed, contextually informed representation. The decoder’s shallow layers employ the LKA [Guo et al.(2023)Guo, Lu, Liu, Cheng, and Hu] modules to adeptly balance local and global features as well as channel interactions, overcoming typical high-resolution image processing challenges associated with self-attention. Deeper decoder layers utilize DAE-Former blocks [Azad et al.(2023)Azad, Arimond, Aghdam, Kazerouni, and Merhof] to efficiently maintain long-range dependencies in lower-resolution images, integrating spatial and channel attentions without disrupting the 2D structure of images. Overall, this method aims at accurate, context-aware segmentation of anatomical structures in complex medical images.
@@ 4.3 Quantitative Results
| ScaleFormer [Huang et al.(2022)Huang, Xie, Lin, Iwamoto, Han, Chen, and Tong] | 111.6 | 48.93 | 89.40 | 83.31 | 86.36 | 74.97 | 95.12 | 80.14 | 88.73 | 64.85 | 82.86 | 16.81 |
@@
| DAEFormer [Azad et al.(2023)Azad, Arimond, Aghdam, Kazerouni, and Merhof] | 48.07 | 27.89 | 91.82 | 82.39 | 87.66 | 71.65 | 95.08 | 80.77 | 87.84 | 63.93 | 82.63 | 16.39 |
@@
| 2D D-LKA Net [Azad et al.(2024)Azad, Niggemeier, Hüttemann, Kazerouni, Aghdam, Velichko, Bagci, and Merhof] | 101.64 | 19.92 | 91.22 | 84.92 | 88.38 | 73.79 | 94.88 | 84.94 | 88.34 | 67.71 | 84.27 | 20.04 |